a particle in a central force field discribed by the spiral orbit r=r0e^(theta). how to show that the force law is inverse cube and that (theta) veries logrithmically with t?.



1
Expert's answer
2021-12-12T16:43:38-0500

"r=r_0e^{b\theta},"

"u=\frac 1r=\frac{e^{-b\theta}}{r_0},"

"\frac{d^2u}{d\theta^2}+u=\frac{F}{mr_0^2u^2},"

"(b^2 \frac{e^{-b\theta}}{r_0}+\frac{e^{-b\theta}}{r_0})\cdot \frac 1{r^2}=\frac F{mr_0^2},"

"(1+b^2)\frac{e^{-b\theta}}{r_0}\cdot \frac 1{r^2}=\frac F{mr_0^2},"

"(1+b^2)\frac 1{r^3}=\frac F{mr_0^2},"

"F\sim \frac 1{r^3}."